OpenStudy (anonymous):

Solve for y: 2/3y + 2 = 8

OpenStudy (aroub):

2/3y+2=8 2/3y*3+2=8*3 y=1/9

OpenStudy (anonymous):

(2/3)y+2=8 (2/3)y+2-2=8-2 (2/3)y=6 (2/3)(3/2)y=6(3/2) 1*y=9 y=9